OpenStudy (anonymous):

What is the geometric mean of 2 and 11

OpenStudy (anonymous):

\[\sqrt{11*2}=\sqrt{22}\]

OpenStudy (anonymous):

Thank you i feel like it was obvious.

OpenStudy (anonymous):

yw, geometric mean of 2 numbers is square root of their product